OpenStudy (anonymous):

ack took 5.5 seconds to fall down the hill. His initial velocity was 0.5 meters per second. His final velocity was 2.5 meters per second. Calculate his average acceleration. I chose the answer 2.0 m/s/s is that correct?

OpenStudy (hoblos):

it is 2m/s/5.5s => it is 0.36m/s/s

OpenStudy (anonymous):

I did that answer last time, but my teacher said that that's incorrect

OpenStudy (anonymous):

I agree with hoblos. Average velocity = \(\Large \frac{{\Delta v}}{{\Delta t}}\)

OpenStudy (anonymous):

Acceleration = Vf - Vi -------- t = (2.5-0.5 )/ 5.5 =0.36 m/s/s I agree as well.

OpenStudy (anonymous):

Sorry, *average acceleration